Question 16 (1 point) 1. Determine the period of the signal with units. V(t) = 80 cos(1000nt - 309) volts.
Answer #1

Given signal V(t)=cos(1000*pi*t-pi/6) volts

Periodic signal: any signal repeats it's shape over a certain time is known as periodic signal. I.e, V(t+T)=v(t). 'T' is the period. T=wo/2*pi.

wo= angular frequency in rad/sec

Given signal is compared with cos(wo*t-pi/6),

Therefore, wo=1000*pi

T=1000*pi/(2*pi)

T=500 seconds

Therefore, Period of the given signal is 500 secs